# Beyond Copenhagen: Following the Trail of Decoherence in Feynman’s Light Microscope

### v2

@@ Abstract
Feynman’s light microscope invites us to reconsider what we thought we knew about quantum reality. Rather than invoking wavefunction collapse to predict the loss of fringes in a monitored interferometer, Feynman analyzes the problem in terms of a disturbance. This approach raises the question of whether the classical world, including its localized particles and definite measurement outcomes, might emerge as the universe evolves smoothly according to Schrödinger’s equation. Treating the particle and its environment as an entangled system, unmodified quantum mechanics shows remarkable success toward this end. This is the purview of decoherence theory. The question of wavefunction collapse then becomes one of what we want from the theory. Do we expect it to describe microscopic reality, or do we consider it to be only a tool for predicting measurement outcomes? Both options are uncomfortable. When applied to unmodified quantum mechanics, the first implies that each moment in time branches into a vast number of divergent macroscopic realities. While the second represents, for most practitioners, a weakened view of science. This article is written to be accessible to anyone with an undergraduate course in quantum mechanics.
@@ I Introduction
Experimental and theoretical developments have gradually changed the situation. Improving technologies allow investigation of quantum behavior of increasingly large systems, calling into question any distinction between microscopic and macroscopic. Molecules with tens of thousands of nuclei have created interference patterns [2], mechanical oscillators on the 100 picogram scale have been prepared in entangled states [3], and mesoscopic fullerene molecules have been shown to decohere smoothly and predictably in the presence of background gas [4] or as they radiate thermal energy [5]. Meanwhile, decoherence theory [6, 7, 8] has demonstrated that environmental entanglement explains how many, if not all, features of classical physics emerge naturally from unbiased quantum mechanics.
Feynman’s light microscope thought experiment [9, 10] can be viewed as a prelude to these developments. An atom scatters a photon while traversing a two-slit interferometer (Fig. 1,) and the photon goes undetected. The Copenhagen theory says that fringe washout occurs if the photon has sufficiently short wavelength to have made possible a which-path measurement, because the atom wavefunction has collapsed into one of the classical single-path states. Feynman had the good sense not to treat such measurement postulates as inscrutable. Since the interaction is a quantum process, might we be able to understand fringe washout without invoking collapse? The light microscope provides a wonderful testing grounds for examining what lies beneath quantum measurement and how the classical world emerges from quantum physics, both theoretically [11, 12, 13, 14, 15, 16, 17, 18] and experimentally [19, 20, 21, 22]. It is remarkable that such a fruitful idea originated from Feynman’s lectures [9] meant for first-year undergraduate students!
Figure 1: Light microscope apparatus.
Feynman’s semiclassical explanation  [^1] Feynman warns that his rough introduction will need revision. Later, in a less well-known portion of the text, he explains fringe washout using entangled states. is that fringe washout occurs because the atom is perturbed by recoil. His semiclassical analysis succeeds at showing the uncertainty principle prevents a logical crisis; it is impossible to observe an atom traveling a single path while also observing interference fringes indicating two-path traversal. However, semiclassical reasoning fails to explain why multiple photons scattered by the single-atom wave will all originate from only one path or the other. It also fails to explain, at a conceptual level, why evidence for classical single-path traversal is more easily obtained than is evidence for non-classical (but equally legitimate) two-path traversal.
@@
Those predictions have so far passed all experimental tests. However, post-Copenhagen quantum mechanics leaves no comfortable option for interpreting reality. Do the equations of physics describe some reality outside our minds? If so, then we are currently experiencing only one of many parallel branches of reality, with most of the other branches being very different from this one. We could instead insist that the only real branch is that corresponding to our experience. But that would necessitate downgrading the wavefunction, from describing any microscopic reality, to being merely a tool for predicting “detector clicks.” Alternatively, we might be dissatisfied with either option and suspect that quantum mechanics needs some fundamental modification, which could change the discussion entirely.
@@ II The Setup
The apparatus (Fig. 1) has a wall in the $z=0$ plane and slits of width $a$, centered at $\pm\frac{d}{2}\hat{x}$. There is an atom detector at $L\hat{z}$. The atom wavepacket approaches the slits from below with central momentum $p_{z}\hat{z}$, and passage through the slits does not change this longitudinal momentum. The diffracted atoms do experience an increase in transverse momentum $p_{x}$, acquired from the wall. Just as it leaves the slits, we can represent the atom in the single-path basis
@@
where $|\text{L}\rangle$ and $|\text{R}\rangle$ describe wavepackets localized at either slit.
@@
Calculation of interference patterns is simpler if the atom detector is in the far-field limit. But then the single-slit diffraction patterns would be indistinguishable from each other and also from the washed-out two-slit pattern. It is worth the trouble to bring the atom detector closer so that those three patterns are distinct, and we use L=10dL=10\,d. Atom patterns are shown in Fig. 2, with the procedure for calculating them described in Appendix A.
The region just above the slits is illuminated with light with wavelength $\lambda$ and $\vec{k}=k\hat{x}$. For simplicity, we consider only $\lambda\gg a$, so that an atom passing through one slit is approximately a point source for scattered photons. The atom recoil from the elastic scattering is $\hbar\vec{\kappa}=\hbar\vec{k}-\hbar\vec{k}_{\text{f}}$, where $\hbar\vec{k}_{\text{f}}$ is the momentum of the scattered photon. Note that the light is not focused, so both sites contribute an amplitude for each scattered photon. (Whether or not those amplitudes are balanced for some detected photon depends on the detection configuration and photon wavelength.)
Figure 3: Photon detection configurations. (a) No detector. (b) Far-field detector. (c) Imaging detector.
@@
We either ignore the scattered photons, or detect them in one of two configurations (Fig. 3.) The first has a photon detector is in the far-field so that each pixel corresponds to a well-defined $\vec{k}_{\text{f}}$. The second configuration uses a lens to image each point $(x,y)$ at $z=0$ onto position $(X=x,Y=y)$ on the detector plane; this is a microscope configuration.
@@ III Anticipating Decoherence
Feynman offers a semiclassical explanation for why fringe washout always accompanies which-path information becoming available. “By trying to ‘watch’ the [atoms] we have changed their motions. That is, the jolt given to the [atom] when the photon is scattered by it is such as to change the [atom’s] motion enough so that if it might have gone to where [the probability] was at a maximum it will instead land where [the probability] was a minimum; that is why we no longer see the wavy interference effects [9].”
@@
Although it is less intuitive, there is a more straightforward way to obtain the same result. To describe the recoiling atom wavefunction, we increase each momentum component of $\psi_{\text{i}}(\vec{r})$ by $\hbar\vec{\kappa}$, such that $\psi(\vec{r})=e^{i\vec{\kappa}\vec{r}}\psi_{\text{i}}(\vec{r})$. Before recoil, the atom wavefunction leaving the left and right slit shared a common phase. After recoil, these phases differ by $\kappa_{x}d$, so the atom fringes locations depend on photon scattering angle. Since $0<\kappa_{x}<2\times 2\pi/\lambda$, averaging over angles washes out fringes if $\lambda\lesssim d$.
Both arguments successfully predict fringe washout when short-wavelength photons are scattered. But by neglecting entanglement, they are limited in two ways. They suggest that decoherence can be blamed on perturbations to the atom as it traverses both paths. If this were the whole story, then with $\lambda\ll d$ and higher light intensity, we should be able observe an atom scattering a first photon from the left path and then a second photon from the right path. The atom wavefunction would not appear to collapse into a position eigenstate after a position measurement, thus contradicting the Copenhagen prediction and experiments [22]. More generally, these semiclassical arguments cannot begin from the underlying quantum mechanics and explain what makes classical behavior special. Why does environmental disturbance seem to steer the atom toward behaving as if it takes a single path, when quantum mechanics should also permit a two-path description?
@@ IV Including Entanglement
Now $|\chi\rangle$ is significantly entangled if $\lambda\lesssim d$, because then the photon resolves the slits, and $\mbox{$|\gamma_{\text{L}}\rangle$}\not\approx\mbox{$|\gamma_{\text{R}}\rangle$}$. For single-slit traversal, if we liked, we could speak separately of an atom state and a scattered photon state. But for two-slit traversal, we have no choice but to speak of a joint atom-photon state.
@@
Since $\mbox{$\langle\text{L},\gamma_{\text{L}},\gamma_{\text{L}}|\chi_{2}\rangle$}\neq 0$, we might see two photons coming from the left path. But $\mbox{$\langle\text{L},\gamma_{\text{R}},\gamma_{\text{L}}|\chi_{2}\rangle$}\propto\mbox{$\langle\gamma_{\text{L}}|\gamma_{\text{R}}\rangle$}$, which vanishes for $\lambda\ll d$ [13]. So, we never see one photon obviously scattered from the left path and the other from the right.
@@ V Different Photon Bases
where $S(\vec{k}_{\text{f}})$ describes the scattering amplitude from an atom located at the origin [13, 15, 16]. The relative phase $\kappa_{x}d=(\vec{k}-\vec{k_{\text{f}}})\cdot d\hat{x}$ can be understood as expressing the same local phase at each slit in terms of the $\vec{k}_{\text{f}}$ mode, offset by the phase accumulated by the incoming photon in traveling between the slits. This interferometer phase shift $\kappa_{x}d$ for each definite-recoil state was anticipated in the semiclassical discussion.
The basis of Eq. 3 is always useful for thinking of photons as originating from definite locations. However, just because the photon has a crisp origination position does not mean it will arrive at a crisp position $X$ on the imaging detector of Fig. 3(c). For considering these images, we use
$\mbox{$|\chi\rangle$}=\frac{1}{\sqrt{2}}\sum_{X}\Bigl{[}\mbox{$\langle X|\gamma_{\text{L}}\rangle$}\mbox{$|\text{L}\rangle$}+\mbox{$\langle X|\gamma_{\text{R}}\rangle$}\mbox{$|\text{R}\rangle$}\Bigr{]}\mbox{$|X\rangle$}.$ (7)
Analytic expressions for the coefficients are given in [13], but we already know from classical optics some of their qualitative features. For $\lambda\ll d$ and a good lens, images of scattering from the two slits will be well-resolved, and $\langle X|\gamma_{\text{L}}\rangle$ will be zero if $\langle X|\gamma_{\text{R}}\rangle$ is non-zero. For $\lambda\gg d$, the images will be completely unresolved and $\mbox{$\langle X|\gamma_{\text{L}}\rangle$}\approx\mbox{$\langle X|\gamma_{\text{R}}\rangle$}$.
@@ VI Decoherence Basics
If the observed atom pattern matches that of two non-interfering sources, as in Fig. 2(b), then each individual atom is said to have decohered. Classical interferometers have various imperfections, such as noisy sources, which can lead to fringe washout. But in the quantum context, the term ‘decoherence’ is generally reserved for that arising from entanglement. The first ingredient for decoherence is environmental entanglement. In the light microscope, the photons are the environment. As we shall see, there are important aspects of (quantum) decoherence which have no classical analogy. The second ingredient for decoherence is ignorance. Decoherence is what happens when we have lost track of some parts an entangled state.
@@
It is now clear that decoherence is intimately tied to entanglement. If $\lambda\ll d$, the scattered photon states are quite different from one another and $\mbox{$\langle\gamma_{\text{R}}|\gamma_{\text{L}}\rangle$}\approx 0$ [13]. In this case, the atom and photon are maximally entangled in Eq. 3, and Eq. 8 predicts complete fringe washout. If $\lambda\gg d$, then $\mbox{$\langle\gamma_{\text{R}}|\gamma_{\text{L}}\rangle$}\approx 1$, $|\chi\rangle$ is nearly factorizable (non-entangled), and we predict survival of interference fringes even though a photon was scattered. The progression of decoherence as the undetected photon wavelength decreases can be seen in the left column of Fig. 4.
Figure 4: Atom probability patterns for different photon detection configurations, with fringes unresolved in the left and right columns. Left: no photon detection. Middle: far-field photon detection. Solid lines show the (arbitrarily scaled) joint atom-photon probability for selected photon scattering angles, corresponding to photons detected at different locations on the detector. Dashed lines show average over all photon angles. Right: imaging detection. Plots show the joint atom-photon probability for photon detection at $X=d/2$ (orange) and at $X=0$ (magenta).
The Born rule by itself predicts that scattering of an undetected short-wavelength photons leads to fringe washout. These photons have potential to provide which-path information to a properly configured detector. There is overlap here with a commonly taught version of the Copenhagen interpretation, which says that the wavefunction collapses if a which-path measurement could potentially be performed, regardless of whether such a measurement is actually performed. Eq. 8 shows that decoherence theory reveals the physics underlying the potentially/actually aspect of this Copenhagen postulate. In any case, there is no need for a macroscopic photon detector for fringe washout to occur.

### v3

@@ Abstract
Feynman’s light microscope invites us to reconsider what we thought we knew about quantum reality. Rather than invoking wavefunction collapse to predict the loss of fringes in a monitored interferometer, Feynman analyzes the problem in terms of a disturbance. This approach raises the question of whether the classical world, including its localized particles and definite measurement outcomes, might emerge as the universe evolves smoothly according to Schrödinger’s equation. Treating the particle and its environment as an entangled system, unmodified quantum mechanics shows remarkable success toward this end. This is the purview of decoherence theory. The question of wavefunction collapse then becomes one of what we want from the theory. Do we expect it to describe microscopic reality, or do we consider it to be only a tool for predicting measurement outcomes? Both options are uncomfortable. The first, when applied to unmodified quantum mechanics, implies that each moment in time branches into a vast number of divergent macroscopic realities. The second represents, for many practitioners, a weakened view of science. This article is written to be accessible to anyone with an undergraduate course in quantum mechanics.
@@ I Introduction
Experimental and theoretical developments have gradually changed the situation. Improving technologies allow investigation of quantum behavior of increasingly large systems, calling into question any distinction between microscopic and macroscopic. Nanoclusters with nearly the mass and diameter of small viruses have created interference patterns [2], mechanical oscillators on the 100 picogram scale have been prepared in entangled states [3], and mesoscopic fullerene molecules have been shown to decohere smoothly and predictably in the presence of background gas [4] and as they radiate thermal energy [5]. Meanwhile, decoherence theory [6, 7, 8] has demonstrated that environmental entanglement explains how many, if not all, features of classical physics emerge naturally from unbiased quantum mechanics.
Feynman’s light microscope thought experiment [9, 10] can be viewed as a prelude to these developments. An atom scatters a photon while traversing a two-slit interferometer (Fig. 1,) and the photon goes undetected. The Copenhagen theory says that we should not see an interference pattern if the photon has sufficiently short wavelength to have made possible a which-path measurement, because the atom wavefunction has collapsed into one of the classical single-path states. Feynman had the good sense not to treat such measurement postulates as inscrutable. Since the interaction is a quantum process, might we be able to understand the disappearance of fringes without invoking collapse? The light microscope provides a wonderful testing grounds for examining, both theoretically [11, 12, 13, 14, 15, 16, 17, 18, 19] and experimentally [20, 21, 22, 23, 24], what lies beneath quantum measurement and how the classical world emerges from quantum physics. It is remarkable that such a fruitful idea originated from Feynman’s lectures [9, 10] meant for first-year undergraduate students!
Figure 1: Light microscope apparatus. The origin for $x$ and $z$ is midway between the slits.
Feynman’s semiclassical explanation  [^1] Feynman warns that his rough introduction will need revision. Later, in a less well-known portion of the text, he explains fringe washout using entangled states. is that fringes average out because the atom is perturbed by recoil. This analysis succeeds at showing the uncertainty principle prevents a logical crisis; it is impossible to observe an atom traveling a single path while also observing interference fringes indicating two-path traversal. However, semiclassical reasoning fails to explain why multiple photons scattered by the single-atom-wave will all originate from only one path or the other. It also fails to explain at a conceptual level what makes classical states special. Why does environmental disturbance seem to steer the atom toward behaving as if it takes a single path, when two-path traversal is equally legal in quantum mechanics?
@@
Those predictions have so far passed all experimental tests. However, post-Copenhagen quantum mechanics leaves no comfortable option for interpreting reality. Do the equations of physics describe some reality outside our minds? If so, then we are currently experiencing only one of many parallel branches of reality, with most other branches looking very different from ours. We could instead insist that the only real branch is this one corresponding to our experience. But to prevent the branching of macroscopic reality, we would need to concede that the wavefunction does not describe any microscopic reality. Rather, it could only be a mathematical tool useful for predicting “detector clicks.” Alternatively, we might be dissatisfied with either option and suspect that quantum mechanics needs some fundamental modification, which could change the discussion entirely.
@@ II The Setup
The apparatus (Fig. 1) has a wall in the $z=0$ plane and slits of width $a$, centered at $\pm\frac{d}{2}\hat{x}$. There is an atom detector at $L\hat{z}$. The atom wavepacket approaches the slits from below with central momentum $p_{0}\hat{z}$, and passage through the slits does not change this longitudinal momentum. The diffracted atom does acquire from the wall a distribution of transverse momenta $p_{x}$. Just as it leaves the slits, we can represent the atom in the single-path basis
@@
where $|\text{L}\rangle$ and $|\text{R}\rangle$ describe wavepackets localized at either slit, and $\mbox{$\langle\text{L}|\text{R}\rangle$}=0$.
@@
The most common configuration is to place the atom detector in the far-field, where calculation of interference patterns is simpler. But then the single-slit diffraction patterns are indistinguishable from each other and also from a decohered two-slit pattern. For our purposes, it is worth the calculational trouble to bring the atom detector closer so that those three patterns are distinct, and we use L=10dL=10\,d. Atom patterns are shown in Fig. 2, and the procedure for calculating them described in Appendix A.
The region just above the slits is illuminated with light of wavelength $\lambda$ and $\vec{k}=k\hat{x}$. For simplicity, we consider only $\lambda\gg a$, so that an atom passing through one slit is approximately a point source for scattered photons. The atom recoil from the elastic scattering is $\hbar\vec{\kappa}=\hbar\vec{k}-\hbar\vec{k}_{\text{f}}$, where $\hbar\vec{k}_{\text{f}}$ is the momentum of the scattered photon. Note that the light is not focused, so both sites contribute an amplitude for each scattered photon. (Whether or not those amplitudes are balanced for any detected photon will turn out to depend on the wavelength and detection configuration.)
@@
Figure 3: Different photon detection configurations (not to scale) to complete Fig. 1. Red wavy arrows represent light scattered from the atom. (a) No photon detection. (b) Far-field photon detection. (c) Imaging detection. (This simplest lens configuration causes images to be inverted, so we define $x_{\gamma}=-x$.)
We either ignore the scattered photons, or detect them in one of two configurations (Fig. 3.) The first has a photon detector in the far-field so that each pixel corresponds to a well-defined $\vec{k}_{\text{f}}$. The second configuration uses a lens to image each position $x$ at $z=0$ onto position $x_{\gamma}=x$ on the detector plane; this is a microscope configuration.
@@ III Anticipating Decoherence
Feynman offers a semiclassical explanation for why fringe washout always accompanies which-path information becoming available. “By trying to ‘watch’ the [atoms] we have changed their motions. That is, the jolt given to the [atom] when the photon is scattered by it is such as to change the [atom’s] motion enough so that if it might have gone to where [the probability] was at a maximum it will instead land where [the probability] was a minimum; that is why we no longer see the wavy interference effects [9, 10].”
@@
Although it is less intuitive, there is a more straightforward way to obtain the same result. To describe the recoiling atom wavefunction, we increase each momentum component of $\psi_{\text{i}}(\vec{r})$ by $\hbar\vec{\kappa}$, such that $\psi(\vec{r})=e^{i\vec{\kappa}\vec{r}}\psi_{\text{i}}(\vec{r})$. Before recoil, the atom wavefunction leaving the left and right slit shared a common phase. After recoil, these phases differ by $\kappa_{x}d$, so the atom fringe locations depend on photon scattering angle. Since $0<\kappa_{x}<2\times 2\pi/\lambda$, averaging over angles washes out fringes if $\lambda\lesssim d$.
Both arguments successfully predict fringe washout when short-wavelength photons are scattered. But by neglecting entanglement, they are limited in two ways. They suggest that decoherence can be blamed on perturbations to the atom as it traverses both paths. If this were the whole story, then with $\lambda\ll d$ and higher light intensity, we should be able observe an atom scattering a first photon from the left path and then a second photon from the right path. The atom wavefunction would not appear to collapse into a position eigenstate after a position measurement, thus contradicting the Copenhagen prediction and experiments [24]. More generally, these semiclassical arguments do not begin with the underlying quantum mechanics and predict emergence of classical behavior.
@@ IV Including Entanglement
Now $|\chi\rangle$ is significantly entangled if $\lambda\lesssim d$, because then the photon resolves the slits, and $\mbox{$|\gamma_{\text{L}}\rangle$}\not\approx\mbox{$|\gamma_{\text{R}}\rangle$}$. For single-slit traversal, we could speak separately of an atom state and a scattered photon state. But for two-slit traversal and $\lambda\lesssim d$, we have no choice but to speak of a joint atom-photon state.
@@
Since $\mbox{$\langle\text{L},\gamma_{\text{L}},\gamma_{\text{L}}|\chi_{2}\rangle$}\neq 0$, we might see two photons coming from the left path. But $\mbox{$\langle\text{L},\gamma_{\text{R}},\gamma_{\text{L}}|\chi_{2}\rangle$}\propto\mbox{$\langle\gamma_{\text{R}}|\gamma_{\text{L}}\rangle$}$, which vanishes for $\lambda\ll d$ [13]. So, we never see one photon obviously scattered from the left path and the other from the right.
@@ V Different Photon Bases
where $S(\vec{k}_{\text{f}})$ describes the scattering amplitude from an atom located at the origin [13, 16, 17]. The relative phase $\kappa_{x}d=(\vec{k}-\vec{k_{\text{f}}})\cdot d\hat{x}$ can be understood as expressing the same local photon phase at each slit in terms of the $\vec{k}_{\text{f}}$ mode, offset by the phase accumulated by the incoming photon in traveling between the slits. This interferometer phase shift $\kappa_{x}d$ for each definite-recoil state was anticipated in the semiclassical discussion.
The basis of Eq. 3 is always useful for treating photons as originating from definite locations. However, just because the photon has a crisp origination position does not mean it will arrive at a crisp position $x_{\gamma}$ on the imaging detector of Fig. 3(c). For considering these images, we use
$\mbox{$|\chi\rangle$}=\frac{1}{\sqrt{2}}\sum_{x_{\gamma}}\Bigl{[}\mbox{$\langle x_{\gamma}|\gamma_{\text{L}}\rangle$}\mbox{$|\text{L}\rangle$}+\mbox{$\langle x_{\gamma}|\gamma_{\text{R}}\rangle$}\mbox{$|\text{R}\rangle$}\Bigr{]}\mbox{$|x_{\gamma}\rangle$}.$ (7)
Analytic expressions for the coefficients are given in [13], but we already know from classical optics some of their qualitative features. For $\lambda\ll d$ and a good lens, images of scattering from the two slits will be well-resolved, and $\langle x_{\gamma}|\gamma_{\text{L}}\rangle$ will be zero if $\langle x_{\gamma}|\gamma_{\text{R}}\rangle$ is non-zero. For $\lambda\gg d$, the images will be completely unresolved and $\mbox{$\langle x_{\gamma}|\gamma_{\text{L}}\rangle$}\approx\mbox{$\langle x_{\gamma}|\gamma_{\text{R}}\rangle$}$.
@@ VI Decoherence Basics
If the atom probability distribution matches that of two non-interfering sources (orange curve in Fig. 2,) then each individual atom is said to have decohered. Classical interferometers have various imperfections, such as spectrally broad sources, which can lead to fringe washout. But in the quantum context, the term ‘decoherence’ is generally reserved for that arising from entanglement. As we shall see, there are important aspects of (quantum) decoherence which have no classical analogy. The first ingredient for decoherence is environmental entanglement. In the light microscope, the photons are the environment. The second ingredient for decoherence is ignorance. Decoherence is what happens when we have lost track of some part of an entangled state.
@@
It is now clear that decoherence is intimately tied to entanglement. If $\lambda\ll d$, the scattered photon states are quite different from one another, and $\mbox{$\langle\gamma_{\text{R}}|\gamma_{\text{L}}\rangle$}\approx 0$ [13]. In this case, the atom and photon are maximally entangled in Eq. 3, and Eq. 8 predicts complete fringe washout. If $\lambda\gg d$, then $\mbox{$\langle\gamma_{\text{R}}|\gamma_{\text{L}}\rangle$}\approx 1$, Eq. 3 is nearly factorizable (non-entangled), and we predict survival of interference fringes even though a photon was scattered. The progression of decoherence, as the undetected photon wavelength decreases, can be seen in the left column of Fig. 4.
Figure 4: Atom probability patterns for different photon detection configurations, with fringes unresolved in the left and right columns. Left: No photon detection. Middle: Far-field photon detection. Solid lines show the atom probability patterns for selected photon scattering angles, corresponding to the photon detected on various pixels. (What is actually plotted is the arbitrarily scaled joint atom-photon probability as a function of $x$, for various $x_{\gamma}$.) Dashed lines show the average over all photon angles. Right: Imaging detection. Plots show the joint atom-photon probability as a function of atom detection at position $x$, for the cases of photon detection at $x_{\gamma}=d/2$ (orange) and at $x_{\gamma}=0$ (magenta).
The Born rule (used in Eq. 8) by itself predicts that scattering of an undetected short-wavelength photons leads to fringe washout. These photons have potential to provide which-path information to a properly configured detector. There is overlap here with a commonly taught version of the Copenhagen theory, which says that the wavefunction collapses if a which-path measurement could potentially be performed, regardless of whether such a measurement is actually performed. Eq. 8 shows that decoherence theory reveals the physics underlying the potentially/actually aspect of this Copenhagen postulate. In any case, there is no need for a macroscopic photon detector for fringe washout to occur.